OpenStudy (anonymous):

What is the relevance of an infinite sum when talking about geometric and arithmetic sequences? (ex. 76 infinity) What and how do you use it?

OpenStudy (amistre64):

If the sequences that is generated approaches zero, then there is a possibility that the sum of the sequence will even out to some solid value ... a1 + a2 + a3 + .... + 0, this is a notion called convergence.

OpenStudy (amistre64):

if the sequence just keep getting bigger then it diverges: a1 + a2 + a3 + ... + inf = inf

OpenStudy (anonymous):

But what if a problem says a geometric sequence u1, u2, u3 ... has u1 = 27 and a sum to infinity os 81/2 what happens?

OpenStudy (amistre64):

well we can generate the infinite formula ... would you agree that the sum of a geometric sequence with a seed of a, and a finite number of terms k a + ar + ar^2 + ar^3 + ar^4 + ... + ar^(k-1) ??

OpenStudy (anonymous):

isnt it a1/1-r

OpenStudy (anonymous):

But what's r?

OpenStudy (amistre64):

well, thats the formula i was going to generate .. so yeah

OpenStudy (amistre64):

do the algebra:\[\frac{81}{2}=\frac{27}{1-r}\] colve for r

OpenStudy (amistre64):

to me it looks like: \[r=1-\frac{27}{81/2}\]

OpenStudy (anonymous):

what does r stand for? like whats the relevance?

OpenStudy (amistre64):

i used it in the last question ... do you remember what i called it and what i used it for?

OpenStudy (amistre64):

since a1 r = a2 ; r = a2/a1 a2 r = a3 ; r = a3/a2 a3 r = a4 ; r = a4/a3 since r is held constant, then there must be a common ratio, r, that relates each new term with the last term used.

OpenStudy (amistre64):

r is called the common ratio and is defined to be \[\large r=\frac{a_{n+1}}{a_n}\]